# wd1: Weighted Policy Optimization for Reasoning in Diffusion Language Models

### v1

@@ 5 Related Work
RL for Diffusion-based LLM. Reinforcement learning for discrete diffusion models has been explored through several approaches. One line of work, exemplified by DRAKES (Wang et al., 2024), leverages reward backpropagation along the denoising trajectory. This approach requires computing a critic and propagating gradients through each denoising step, which is computationally intensive and prone to vanishing gradients. Alternatively, methods such as MMaDA (Yang et al., 2025) and d1 (Zhao et al., 2025) adopt direct RL formulations like GRPO, approximating missing diffusion components—such as per-token likelihoods—for policy optimization. Zhu et al. (2025a) applies Direct Preference Optimization (DPO) to fine-tune the LLaDA base model (Nie et al., 2025a), achieving notable gains in reasoning tasks. However, these approaches all depend on likelihood ratios, which can introduce bias and instability due to likelihood approximation errors. In contrast, our method derives a weighted policy optimization approach that eliminates the need for explicit policy ratios. Importantly, similar to prior works, our method directly optimizes the predictive distribution over clean data. A complementary line of research formulates policy optimization in terms of concrete scores (Lou et al., 2024; Meng et al., 2022). SEPO (Zekri and Boullé, 2025), for instance, introduces a policy optimization objective that only depends on concrete score estimation, thereby circumventing likelihood approximation altogether.
@@ 6 Conclusion
In this work, we introduce wd1, a weighted policy optimization method for reasoning with diffusion-based large language models (dLLMs). wd1 is designed to minimize reliance on likelihood approximation in policy optimization, thereby mitigating the potentially substantial bias that can arise from approximation errors in policy ratios. Our method is grounded in a weighted log-likelihood objective, derived to approximate the closed-form solution to the reverse-KL-constrained policy optimization. Empirically, we show that wd1, even without supervised fine-tuning, surpasses the existing method d1 by up to 16% in accuracy on reasoning benchmarks, while also delivering notable improvements in computational efficiency during RL training. These results highlight the effectiveness of wd1 and establish it as a more scalable and efficient approach for fine-tuning dLLMs.
Appendix A Theoretical Insights
@@
Theorem 1 (Policy Improvement Bound (Kakade and Langford, 2002; Schulman et al., 2015)).
@@
$\eta(\pi_{\theta^{*}})\geq L_{\pi_{\text{old}}}(\pi_{\theta^{*}})-CD_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{TV}}}^{\max}(\pi_{\text{old}},\pi_{\theta^{*}})^{2}.$
@@
Based on Theorem 1, due to $D_{\text{TV}}^{\max}(a||b)^{2}\leq D_{\text{KL}}^{\max}(a||b)$ (Pollard, 2000; Schulman et al., 2015), TRPO and PPO with fixed FKL regularization have the monotonic improvement guarantees. In other words, $\eta(\pi_{\theta^{*}})\geq L_{\pi_{\text{old}}}(\pi_{\theta^{*}})-CD_{{\text{TV}}}^{\max}(\pi_{\text{old}},\pi_{\theta^{*}})^{2}\geq L_{\pi_{\text{old}}}(\pi_{\theta^{*}})-C\mathbb{E}[D_{\text{KL}}(\pi_{\text{old}}\|\pi_{\theta^{*}})]\geq L_{\pi_{\text{old}}}(\pi_{\text{old}})=\eta(\pi_{\text{old}})$.
Theorem 2.
The NLL objective satisfies that
$\mathcal{L}_{\text{NLL}}:=-\mathbb{E}_{q\sim\mathcal{D},o\sim\pi_{\theta^{*}}}[\log\pi_{\theta}(o|q)]\approx-\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(\cdot|q)}\Big{[}\sum_{i=1}^{\mathrm{G}}\frac{\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o_{i})\big{)}}{\sum_{j=1}^{\mathrm{G}}\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o_{j})\big{)}}\cdot\log\pi_{\theta}(o_{i}|q)\Big{]}.$ (10)
@@
To obtain the practical objective in Equation 7, we first start from the cross-entropy loss, and obtain the following:
$\mathcal{L}_{\text{NLL}}$ $=-\mathbb{E}_{q\sim\mathcal{D},\ o\sim\pi_{\theta^{*}}(\cdot|q)}\Big{[}\log\pi_{\theta}(o|q)\Big{]}$ (11)
$=-\mathbb{E}_{q\sim\mathcal{D}}\Big{[}\sum_{o}\pi_{\theta^{*}}(o|q)\cdot\log\pi_{\theta}(o|q)\Big{]}$ (12)
$=-\mathbb{E}_{q\sim\mathcal{D}}\Big{[}\sum_{o}\frac{\pi_{\text{old}}^{\text{ref}}(o|q)\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o)\big{)}}{\sum_{o^{\prime}}\pi_{\text{old}}^{\text{ref}}(o^{\prime}|q)[\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o^{\prime})\big{)}]}\cdot\log\pi_{\theta}(o|q)\Big{]}$ (13)
$=-\mathbb{E}_{q\sim\mathcal{D},o\sim\pi_{\text{old}}^{\text{ref}}(o|q)}\Big{[}\frac{\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o)\big{)}}{\mathbb{E}_{o\sim\pi_{\text{old}}^{\text{ref}}}[\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o)\big{)}]}\cdot\log\pi_{\theta}(o|q)\Big{]}$ (14)
$\approx-\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(o|q)}\Big{[}\frac{1}{G}\sum_{i=1}^{\mathrm{G}}\frac{\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o_{i})\big{)}}{\mathbb{E}_{o\sim\pi_{\text{old}}^{\text{ref}}}[\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o)\big{)}]}\cdot\log\pi_{\theta}(o_{i}|q)\Big{]}.$ (15)
In the above we went from Equation 12 to Equation 13 by simply using the know form of the optimal policy $\pi^{*}(\cdot|q)\propto\pi_{{\text{old}}}^{\text{ref}}(\cdot|q)\cdot\exp\big{(}\psi\hat{A}(q,\cdot)\big{)}$. We go from Equation 13 to (14) by using the definition of expectation and from Equation 14 to Equation 15 by approximating through $G$ samples $\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(o|q)$.
Finally, we can approximate the normalization constant $\mathbb{E}_{o\sim\pi_{\text{old}}^{\text{ref}}}\big{[}\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o)\big{)}\big{]}$ using sampling, thus:
$\mathcal{L}_{\text{NLL}}$ $\approx-\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(\cdot|q)}\Big{[}\frac{1}{G}\sum_{i=1}^{\mathrm{G}}\frac{\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o_{i})\big{)}}{\frac{1}{G}\sum_{j=1}^{\mathrm{G}}\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o_{j})\big{)}}\cdot\log\pi_{\theta}(o_{i}|q)\Big{]}$ (16)
$=-\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(\cdot|q)}\Big{[}\sum_{i=1}^{\mathrm{G}}w(q,o_{i})\cdot\log\pi_{\theta}(o_{i}|q)\Big{]},$ (17)
where $w(q,o_{i})=\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o_{i})\big{)}/\sum_{j=1}^{\mathrm{G}}\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o_{j})\big{)}$. ∎
@@
then $\eta^{\prime}(\pi_{\theta^{*}})\geq\eta^{\prime}(\pi_{\text{old}})$.
@@
Based on Theorem 1, we have
$\eta^{\prime}(\pi_{\theta^{*}})=$ $\eta(\pi_{\theta^{*}})-\mathbb{E}_{q\in\mathcal{D}}\Big{[}\beta D_{\text{KL}}\Big{(}\pi_{\theta^{*}}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big{)}\Big{]}$
$\geq$ $L(\pi_{\theta^{*}})-CD_{\text{TV}}^{\max}(\pi_{\text{old}},\pi_{\theta^{*}})^{2}-\mathbb{E}_{q\in\mathcal{D}}\Big{[}\beta D_{\text{KL}}\Big{(}\pi_{\theta^{*}}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big{)}\Big{]}$ (20)
$\geq$ $L(\pi_{\theta^{*}})-CD_{\text{KL}}^{\max}(\pi_{\theta^{*}}\|\pi_{\text{old}})-\mathbb{E}_{q\in\mathcal{D}}\Big{[}\beta D_{\text{KL}}\Big{(}\pi_{\theta^{*}}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big{)}\Big{]}$ (21)
$\geq$ $L(\pi_{\theta^{*}})-\mathbb{E}_{q\in\mathcal{D}}\Big{[}\lambda D_{\text{KL}}(\pi_{\theta^{*}}(\cdot|q)\|\pi_{\text{old}}(\cdot|q))+\beta D_{\text{KL}}\Big{(}\pi_{\theta^{*}}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big{)}\Big{]}$ (22)
$=$ $M(\pi_{\theta^{*}})$ (23)

### v2

@@ 6 Related Work
RL for Diffusion-based LLM. RL for discrete diffusion models has been explored through several approaches. One line of work, exemplified by DRAKES (Wang et al., 2024), leverages reward backpropagation along the denoising trajectory. This approach requires computing a critic and propagating gradients through each denoising step, which is computationally intensive and prone to vanishing gradients. Alternatively, methods such as MMaDA (Yang et al., 2025) and d1 (Zhao et al., 2025) adopt direct RL formulations like GRPO, approximating missing diffusion components—such as per-token likelihoods—for policy optimization. Zhu et al. (2025a) applies Direct Preference Optimization (DPO) to fine-tune the LLaDA base model (Nie et al., 2025a), achieving notable gains in reasoning tasks. However, these approaches all depend on likelihood ratios, which can introduce bias and instability due to likelihood approximation errors. In contrast, our method derives a weighted policy optimization approach that eliminates the need for explicit policy ratios. Importantly, similar to prior works, our method directly optimizes the predictive distribution over clean data. A complementary line of research formulates policy optimization in terms of concrete scores (Lou et al., 2024; Meng et al., 2022). SEPO (Zekri and Boullé, 2025), for instance, introduces a policy optimization objective that only depends on concrete score estimation, thereby circumventing likelihood approximation altogether.
@@
"Ratio-Free" Policy Optimization. If a policy optimization objective requires neither importance sampling nor regularization with respect to a reference model, then the objective is ratio-free. Consequently, on-policy algorithms such as vanilla policy gradient methods (e.g., REINFORCE (Williams, 1992)) and their variants (e.g., RLOO (Kool et al., 2019)) are inherently ratio-free. This property is particularly valuable for dLLMs, where errors in log-likelihood approximation can accumulate and propagate through ratio-based computations. Concurrent work, such as SPG (Wang et al., 2025a), adopts a policy-gradient formulation and develops an objective tailored specifically for diffusion language models. Another on-policy optimization approach, d2 (Wang et al., 2025c), removes both the ratios and the likelihood terms from the RL objective for dLLMs, offering a more fundamental solution. However, our method wd1, similar to AWR (Peng et al., 2019), is inherently an off-policy loss, which is more general.
@@ 7 Conclusion
We introduce wd1, a weighted policy optimization method for reasoning with dLLMs. wd1 is designed to minimize reliance on likelihood approximation, thereby mitigating the potentially substantial bias that can arise from approximation errors in policy ratios. Our method is grounded in a weighted log-likelihood objective, derived to approximate the closed-form solution to the reverse-KL-constrained policy optimization. Empirically, we show that wd1, even without supervised fine-tuning, surpasses the existing method d1 by up to 16% in accuracy on reasoning benchmarks, while also delivering notable improvements in computational efficiency during RL training. These results highlight the effectiveness of wd1 and establish it as a more scalable and efficient approach for fine-tuning dLLMs.
8 Ethics and Reproducibility Statement
This work raises no question or concern regarding the Code of Ethics. As for reproducibility of our results, we provide details of implementations in Section˜5, in Experimental Setup and Implementation subsections. Additional details including dataset, reward functions, and hyperparameters are provided in Appendix˜B. All the theoretical results are proved in Appendix˜A.
9 Acknowledgments
Ilija Bogunovic was supported by the EPSRC New Investigator Award EP/X03917X/1; the Engineering and Physical Sciences Research Council EP/S021566/1. Sangwoong Yoon was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2020-II201336, Artificial Intelligence Graduate School Program (UNIST)), the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2024-00408003), and the Center for Advanced Computation at Korea Institute for Advanced Study. Xiaohang Tang was supported by the Engineering and Physical Sciences Research Council [grant number EP/T517793/1, EP/W524335/1]. Rares Dolga is supported by EPSRC, grant reference number EP/S021566/1.
The authors would like to thank UIPath and Che Liu (Imperial College London) for providing computing resources that supported our experiments, and Prof. David Barber, Yiming Yang, Xiaoyuan Cheng, and Keyue Jiang from University College London for valuable discussions during the early stages of this work.
Appendix A Proofs and Additional Theory
A.1 Objective Estimation Error due to Likelihood Approximation
In this section, we aim to show that diffu-GRPO amplify the log-likelihood approximation error. Denote the approximator by $\phi$ such that $\|\phi^{\pi_{\theta}}(q,o)-\log\pi_{\theta}(o|q)\|\leq\epsilon$ and $\|\phi^{\pi_{\text{old}}}(q,o)-\log\pi_{\text{old}}(o|q)\|\leq\epsilon^{\prime}$. Then the objective diffu-GRPO in the worst case suffers from exponential error. We discuss the case without ratio clipping and omit the regularization for convenience. Denote $\mathcal{L}_{\text{GRPO}}$ as the ground truth objective without likelihood approximation:
$\|$ $\mathcal{L}_{\text{{diffu}-GRPO}}-\mathcal{L}_{\text{GRPO}}\|$
$=$ $||\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\big(\exp{\phi^{\pi_{\theta}}(o_{i}^{k})}/\exp{\phi^{\pi_{\text{old}}}(o_{i}^{k})}\big)\hat{A}_{i}\big]\Bigg]$
- $\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\big(\pi_{\theta}(o_{i}^{k})/\pi_{\text{old}}(o_{i}^{k})\big)\hat{A}_{i}\Bigg]||$
$=$ $||\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\exp{\big(\phi^{\pi_{\theta}}(o_{i}^{k})-\phi^{\pi_{\text{old}}}(o_{i}^{k})\big)}\hat{A}_{i}\big]\Bigg]$
- $\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\big(\pi_{\theta}(o_{i}^{k})/\pi_{\text{old}}(o_{i}^{k})\big)\hat{A}_{i}\big]\Bigg]||$
$\leq$ $||\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\exp{\big(\log\pi_{\theta}(o_{i}^{k})-\log\pi_{\text{old}}(o_{i}^{k})+(\epsilon+\epsilon^{\prime})\big)}\hat{A}_{i}\big]\Bigg]$
- $\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\big(\pi_{\theta}(o_{i}^{k})/\pi_{\text{old}}(o_{i}^{k})\big)\hat{A}_{i}\big]\Bigg]||$
$=$ $||\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\exp{\big(\epsilon+\epsilon^{\prime}\big)}\hat{A}_{i}\big]\Bigg]\leq C\exp{\big(\epsilon+\epsilon^{\prime}\big)},$ (15)
where $C$ is a constant independent to $\epsilon$ and $\epsilon^{\prime}$. In contrast wd1 has only linear approximation error. Denote the objective computed using approximated log-likelihood as $\mathcal{L}_{\phi}$
$\|\mathcal{L}_{\phi}-\mathcal{L}_{\textit{wd1}}\|=$ $\|\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}^{\text{ref}}(\cdot|q)}\Big[\sum_{i=1}^{G}\big(-w^{+}(q,o_{i})+w^{-}(q,o_{i})\big)\cdot\big(\phi(q,o_{i})-\log\pi_{\theta}(o_{i}|q)\big)\Big]\|$
$\leq$ $C^{\prime}\epsilon.$ (16)
A.2 Reinforcement Learning
@@
Proposition 1 (Policy Improvement Bound (Kakade and Langford, 2002; Schulman et al., 2015)).
@@
$\eta(\pi^{*})\geq L_{\pi_{\text{old}}}(\pi^{*})-CD_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{TV}}}^{\max}(\pi_{\text{old}},\pi^{*})^{2}.$
@@
Based on Proposition˜1, due to $D_{\text{TV}}^{\max}(a||b)^{2}\leq D_{\text{KL}}^{\max}(a||b)$ (Pollard, 2000; Schulman et al., 2015), TRPO and PPO with fixed forward KL regularization have the monotonic improvement guarantees. In other words, $\eta(\pi^{*})\geq L_{\pi_{\text{old}}}(\pi^{*})-CD_{{\text{TV}}}^{\max}(\pi_{\text{old}},\pi^{*})^{2}\geq L_{\pi_{\text{old}}}(\pi^{*})-C\mathbb{E}[D_{\text{KL}}(\pi_{\text{old}}\|\pi^{*})]\geq L_{\pi_{\text{old}}}(\pi_{\text{old}})=\eta(\pi_{\text{old}})$.
Proposition 2.
Minimizing $D_{\text{KL}}(\pi^{*}(\cdot|q)\,\|\,\pi_{\theta}(\cdot|q))$ w.r.t. $\theta$ is equivalent to optimize the following loss objective:
$\mathcal{L}_{\text{WLL}}(\theta)=$ $\mathbb{E}_{q\sim\mathcal{D},o\sim\pi^{\text{ref}}_{\text{old}}(\cdot|q)}\Big[-\exp\big(\psi A^{\pi_{\text{old}}}(q,o)\big)\cdot\log\pi_{\theta}(o_{i}|q)\Big]$ (17)
$\approx$ $-\mathbb{E}_{\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(o|q)}\Big[\frac{1}{G}\sum_{i=1}^{\mathrm{G}}\frac{\exp\big(\psi A^{\pi_{\text{old}}}(q,o_{i})\big)}{\sum_{j=1}^{\mathrm{G}}[\exp\big(\psi A^{\pi_{\text{old}}}(q,o_{j})\big)]}\cdot\log\pi_{\theta}(o_{i}|q)\Big].$ (18)
@@
To obtain the practical objective in Equation˜18, we first start from the cross-entropy loss, and obtain the following. $\forall q\in\mathcal{D}$:
$D_{\text{KL}}(\pi^{*}(\cdot|q)\,\|\,\pi_{\theta}(\cdot|q))$
$=-\mathbb{E}_{o\sim\pi^{*}(\cdot|q)}\Big[\log\pi_{\theta}(o|q)\Big]$ (19)
$=-\Big[\sum_{o}\pi^{*}(o|q)\cdot\log\pi_{\theta}(o|q)\Big]$ (20)
$=-\Big[\sum_{o}\frac{\pi_{\text{old}}^{\text{ref}}(o|q)\exp\big(\psi A^{\pi_{\text{old}}}(q,o)\big)}{\sum_{o^{\prime}}\pi_{\text{old}}^{\text{ref}}(o^{\prime}|q)[\exp\big(\psi A^{\pi_{\text{old}}}(q,o^{\prime})\big)]}\cdot\log\pi_{\theta}(o|q)\Big]$ (21)
$=-\mathbb{E}_{o\sim\pi_{\text{old}}^{\text{ref}}(o|q)}\Big[\frac{\exp\big(\psi A^{\pi_{\text{old}}}(q,o)\big)}{\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}^{\text{ref}}}[\exp\big(\psi A^{\pi_{\text{old}}}(q,o^{\prime})\big)]}\cdot\log\pi_{\theta}(o|q)\Big]$ (22)
Since the normalization constant $\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}^{\text{ref}}}[\exp\big(\psi A^{\pi_{\text{old}}}(q,o^{\prime})\big)]$ is independent to $o$, we can convert the objective to a weighted log-likelihood, and approximate it with samples from the group and weight normalization to obtain:
$\mathcal{L}_{\text{WLL}}(\theta)=$ $-\mathbb{E}_{o\sim\pi_{\text{old}}^{\text{ref}}(o|q)}\Big[{\exp\big(\psi A^{\pi_{\text{old}}}(q,o)\big)}\cdot\log\pi_{\theta}(o|q)\Big]$ (23)
$\approx$ $-\mathbb{E}_{\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(o|q)}\Big[\frac{1}{G}\sum_{i=1}^{\mathrm{G}}\frac{\exp\big(\psi A^{\pi_{\text{old}}}(q,o_{i})\big)}{\sum_{j=1}^{\mathrm{G}}[\exp\big(\psi A^{\pi_{\text{old}}}(q,o_{j})\big)]}\cdot\log\pi_{\theta}(o_{i}|q)\Big].$ (24)
We derive Equation˜21 from Equation˜20 by simply using the known form of the optimal policy $\pi^{*}(\cdot|q)\propto\pi_{{\text{old}}}^{\text{ref}}(\cdot|q)\cdot\exp\big(\psi\hat{A}(q,\cdot)\big)$. We derive Equation˜22 from Equation˜21 by using the definition of expectation and from Equation˜22 to Equation˜24 by approximating through $G$ samples $\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(o|q)$.
@@
then $\eta^{\prime}(\pi^{*})\geq\eta^{\prime}(\pi_{\text{old}})$.
@@
Based on Proposition˜1, we have
$\eta^{\prime}(\pi^{*})=$ $\eta(\pi^{*})-\mathbb{E}_{q\in\mathcal{D}}\Big[\beta D_{\text{KL}}\Big(\pi^{*}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$
$\geq$ $L(\pi^{*})-CD_{\text{TV}}^{\max}(\pi_{\text{old}},\pi^{*})^{2}-\mathbb{E}_{q\in\mathcal{D}}\Big[\beta D_{\text{KL}}\Big(\pi^{*}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$ (27)
$\geq$ $L(\pi^{*})-CD_{\text{KL}}^{\max}(\pi^{*}\|\pi_{\text{old}})-\mathbb{E}_{q\in\mathcal{D}}\Big[\beta D_{\text{KL}}\Big(\pi^{*}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$ (28)
$\geq$ $L(\pi^{*})-\mathbb{E}_{q\in\mathcal{D}}\Big[\lambda D_{\text{KL}}(\pi^{*}(\cdot|q)\|\pi_{\text{old}}(\cdot|q))+\beta D_{\text{KL}}\Big(\pi^{*}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$ (29)
$=$ $M(\pi^{*})$ (30)